# Optimal Transport Aggregation for Visual Place Recognition

### v1

@@ 1 Introduction
Regarding aggregation, the handcrafted VLAD [26] and its learned counterpart NetVLAD [4] are among the most popular choices. Basically, they aggregate local descriptors by quantizing them into a set of clusters and storing the sum of residuals per cluster. Alternative methods include pooling layers like GeM [44] or learned global aggregation, like the recent MixVPR [2]. In this paper, we propose optimal transport aggregation, setting a new state of the art in VPR.
@@
  • Secondly, we integrate the representational power of foundation models into VPR, using DINOv2 model as the backbone for feature extraction. Unlike previous approaches that utilized DINOv2 in its pre-trained form, our method involves fine-tuning the model specifically for the task. This fine-tuning process converges extremely fast, in just four epochs, and allows DINOv2 to capture more relevant and distinctive features pertinent to place recognition tasks.
The fusion of these two novel components results in DINOv2 SALAD, which can be efficiently trained in less than one hour and sets unprecedented recalls in VPR benchmarks, with 75.0% Recall@1 in MSLS Challenge and 76.0% in Nordland. All of this with a single-stage pipeline, without requiring expensive post-processing steps and with an inference speed of less than 3 ms per image.
@@ 3.1 Local Feature Extraction
DINOv2 adopts a ViT architecture that initially divides an input image $\mathbf{I}\in\mathbb{R}^{h\times w\times c}$ into $p\times p$ patches, with $p=14$. These patches are sequentially projected with transformer blocks, resulting in the output tokens $\{\mathbf{t}_{1},\dots,\mathbf{t}_{n},\mathbf{t}_{n+1}\},\mathbf{t}_{i}\in\mathbb{R}^{d}$, where $n=hw/p^{2}$ is the number of input patches and there is an added global token $\mathbf{t}_{n+1}$ that aggregates class information. Although the DINOv2’s authors reported that fine-tuning the model only brings dim improvements, we found that at least for VPR there are substantial gains in selectively unfreezing and training the last blocks of the encoder.
@@ 3.3 Aggregation
where $\textbf{t}_{n+1}$ is the global token from DINOv2. We then concatenate g with V flattened. Following NetVLAD, we do an L2 intra-normalization and an entire L2 normalization of this vector, yielding the final global descriptor.
@@ 4 Experiments
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
NetVLAD [4] 35.1 47.4 51.7 82.6 89.6 92.0 32.6 47.1 53.3 90.5 96.2 97.4 78.7 88.3 91.4
GeM [44] 49.7 64.2 67.0 78.2 86.6 89.6 21.6 37.3 44.2 87.0 94.4 96.3 66.7 83.4 88.0
CosPlace [5] 60.9 71.7 76.7 85.4 90.7 92.3 43.8 60.1 66.8 92.4 97.2 98.1 79.6 90.4 92.8
MixVPR [2] 64.0 75.9 80.6 88.0 92.7 94.6 58.4 74.6 80.0 94.6 98.3 99.0 85.2 92.1 94.6
EigenPlaces [6] 67.4 77.1 81.7 89.3 93.7 95.0 54.4 68.8 74.1 94.1 98.0 98.7 69.9 82.9 87.6
DINOv2 SALAD (ours) 75.0 88.8 91.3 92.2 96.4 97.0 76.0 89.2 92.0 95.1 98.5 99.1 92.1 96.2 96.5
Table 1: Comparison against single-stage baselines. We compare our DINOv2 SALAD against two popular baselines [4, 44] and the three baselines that show best results in recent literature [5, 2, 6]. Our results are the state of the art in all metrics, in most cases by a significant margin. Note, in particular, the large improvement in the most challenging benchmarks, MSLS Challenge and NordLand. † We reproduced GeM and CosPlace results training during 80 epochs following MixVPR training pipeline .
Method Feature Dim Memory (GB) Latency (ms) MSLS Challenge MSLS Val
@@
Table 2: Comparison against baselines with re-ranking. We compare our single-stage DINOv2 SALAD with methods that perform a re-ranking stage to improve performance. Without using re-ranking, our DINOv2 SALAD outperforms all other methods while being orders of magnitude faster. Latency metrics obtained from [64] using a RTX A5000 GPU. Latency for DINOv2 SALAD was computed using a RTX 3090. Memory footprint is calculated on the MSLS Val dataset, which includes around $18,000$ images.
To rigorously evaluate the effectiveness of our proposed contributions, we conducted exhaustive experiments following standard evaluation protocols. In Sec. 4.1, we present implementation details regarding architecture, training, and validation. Then, in  Sec. 4.2, we compare our method with recent VPR methods, and, in  Sec. 4.3, we show ablation studies that assess the importance of the different proposed contributions. Section 4.4 shows qualitative results that help to interpret SALAD’s components.
@@ 4.1 Implementation Details
For the architecture, we opt for a pretrained DINOv2-B as our feature extraction backbone, targeting a balance between computational efficiency and representational capacity. We keep frozen most of the backbone with exception of the the final 4 layers of the encoder. This approach enhances performance significantly without markedly increasing training time. For the fully connected layers, the weights of the hidden layers $\mathbf{W}_{s_{1}}$, $\mathbf{W}_{f_{1}}$ and $\mathbf{W}_{g_{1}}$ have $512$ neurons and use ReLU for the activation function $\sigma$. To optimize feature handling, we employ a dimensionality reduction, compressing feature and global token dimensions from $d=768$ to $l=128$. We use $m=64$ clusters, resulting in a global descriptor of size $128\times 64+256$.
We train on GSV-Cities [1], a large dataset of urban locations collected from Google Street View. Given the impressive representation power of DINOv2, our pipeline achieves training convergence within just 4 complete epochs. Using a batch size of 60 places, each represented by 4 images, the training is completed in 30 minutes on a single NVIDIA 3090GTX. We use Multi-Similarity loss [58] and AdamW [34] for the optimization, with a learning rate set to $6\mathrm{e}{-5}$. To ensure an effective learning rate, we reduce the initial rate linearly at every step until $20\%$ of the initial value. We use dropout of $0.3$ on the score projection and dimensionality reduction neurons. As our model is agnostic to the image input size (as long as it can be divided in $14\times 14$ patches), we evaluate on images of size $322\times 322$ but train on $224\times 224$ to speedup training time.
To validate our experiments and obtain the best set of hyperparameters, we monitored the recall in the Pittsburg30k-test [56]. We observed that in the long run, most configurations perform similarly, but rapid convergence on a few epochs is more sensitive to the hyperparameters.
@@ 4.2 Results
We benchmarked our model against several single-stage baselines, namely NetVLAD [4] and GeM [44] as two representative tradicional baselines, and CosPlace [5], MixVPR [2] and EigenPlaces [6] as the three most recent and best performing baselines in the literature. The evaluation spanned a diverse array of well-established datasets: MSLS Validation and Challenge [59], which are comprised of dashcam images; Pittsburgh30k-test and Pittsburgh250k-test [56], featuring urban scenarios; SPED [14], a collection from surveillance cameras; and NordLand, notable for its seasonal variations from images captured from the front of a train traversing Norway. We use Recall@k (R@k) as the metric for all our experiments, as it is standard in related work. We use evaluation data and code from MixVPR [2], which considers retrieval as correct if an image at less than $25$ meters from the query is among the top-k predicted candidates.
As shown in Table 1, our model outperforms all previous methods on all datasets and all metrics. It is worth highlighting the metrics saturation observed in MSLS Val, Pitts250k-test and SPED, and on the other hand the challenging nature of MSLS Challenge and NordLand, for which all baselines show lower R@k. The MSLS Challenge dataset, with its diversity, extensive size and closed labels, and NordLand, with its extreme sample similarity and seasonal shifts, emerge then as key benchmarks for assessing VPR performance. Although our DINOv2 SALAD shows a significant improvement on all benchmarks, it is precisely in MSLS Challenge and NordLand where we obtain the most substantial recall increases, with $+7.6\%,+11.7\%,+9.6\%$ and $+17.6\%,+14.6\%,+12.0\%$ for R@1,R@5,R@10 respectively over the second best.
In Table 2, we compare our DINOv2 SALAD method, which solely operates on a single retrieval stage, against the leading two-stage Visual Place Recognition (VPR) techniques. In this comparison, we include the best performing models in the literature, namely R2Former [64], TransVPR [57], and Patch-NetVLAD [23], which incorporate a re-ranking refinement. Note in this table how our DINOv2 SALAD, despite being orders of magnitude faster and smaller in memory, significantly outperforms all these two-stage methods on all benchmarks. This finding not only highlights the efficiency of our model but also demonstrates the effectiveness of global retrieval using our novel SALAD aggregation. Additionally, considering our method’s reliance on local features, we believe that a re-ranking stage could also be applied on top of these, potentially increasing even more our recall metrics but at the price of a higher computational footprint.
MSLS Challenge MSLS Val NordLand Pitts250k-test SPED
Feature Dim R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
@@ 4.3 Ablation Studies
In this section, we present different studies that evaluated the efficacy of the components and configurations in our proposed methods.
@@
Effect of DINOv2. We assess the impact of the DINOv2 backbone and our optimal transport aggregation SALAD separately. For this, we compare with the existing baselines of ResNet NetVLAD or AnyLoc, this last one applying a VLAD on top of a pretrained DINOv2 encoder. We integrate the DINOv2 backbone with various aggregation modules, obtaining a handful of performant techniques that improve their respective previous results. As shown in Table 3, all of these configurations outperform the baselines, even though AnyLoc already uses DINOv2. This validates the DINOv2’s integration in end-to-end fine-tuning to refine its feature extraction capabilities.
@@
Effect of SALAD. Our experiments in Table 3 show that aggregation also matters. Even the recent MixVPR aggregation coupled with DINOv2 does not match the performance of DINOv2 NetVLAD and DINOv2 SALAD. We believe that the DINOv2 backbone is especially suitable for local feature aggregation, as its features work remarkably well in dense visual perception tasks [40, 27, 61]. Although DINOv2 NetVLAD achieves comparable performance to SALAD, it employs a descriptor almost three times as big. Besides, the generalization performance of DINOv2 NetVLAD is limited, as observed in NordLand results. We attribute this to NetVLAD’s priors initialization with urban scenarios, which constrain the convergence of the system. In our experiments we also trained a slimmer DINOv2 NetVLAD version, whose features are dimensionally reduced as described in Section 3.3, targetting a final descriptor of roughly the same size as SALAD. In this fairer setup, DINOv2 SALAD clearly outperforms DINOv2 NetVLAD.
Figure 3: Recall@1 in MSLS Val for different configurations of the DINOv2 backbone and the dimension of the aggregated features, $l$.
@@
Effect of hyperparameters. DINOv2 comes in different sizes (Table 4) that affect the number of parameters, inference speed, and representation capabilities. As shown in Figure 3, more parameters do not always result in better performance. Excesively big models might be harder to train or prone to overfit the training set. From these results, we chose the DINOv2-B backbone. Regarding the dimension reduction, we observed that the Recall@1 is quite stable for different descriptor sizes, with a slight peak for $128$ dimensions and worse performance after that. A similar trade-off arises in Table 6 for the number of blocks to train. We observed that fine-tuning two or four blocks report the best results without significant computation overhead.
Model Dim. size Num. blocks Num. parameters
S 384 12 21M
B 768 12 86M
L 1024 24 300M
G 1536 40 1100M
Table 4: DINOv2 configurations. We chose DINOv2-B for its balance between capacity and size.
@@
Figure 4: Heatmap of local features importance. Left images show the original pictures, their right counterparts represent the weights not assigned to the ‘dustbin’. Note how the network learns to discard uninformative regions like skies, roads or dynamic objects, and instead focus on distinctive patterns in buildings and vegetation.
@@
Effect of SALAD components. In Table 5, we show how different components of our SALAD pipeline affect the final performance. Both the global token, which appends global information not captured in local features, and the dustbin, which helps distill the aggregated features, contribute to the performance of SALAD. We also trained a model using a dual-softmax [45] to solve the optimal transport assignment, following LoFTR and Gluestick [51, 41]. Although it achieves only slightly worse performance, the Sinkhorn Algorithm its theoretically sound and provides a better acronym to our method.
Method MSLS Val
R@1 R@5 R@10
DINOv2 SALAD w/o dustbin 91.4 95.8 96.2
DINOv2 SALAD w/o global token 91.8 96.0 96.2
DINOv2 SALAD (Dual Softmax) 91.9 95.7 96.5
DINOv2 SALAD 92.2 96.4 97.0
Table 5: Ablation study of the SALAD components. All components play a role in our state-of-the-art results. In general, the dustbin component is the most influential in the recall, then the global token, and finally the optimal transport method chosen.
@@
Table 6: Effect of fine-tuning different number of DINOv2 blocks.

### v2

@@ 1 Introduction
Regarding aggregation, NetVLAD [4], the learned counterpart to the traditional handcrafted VLAD [26], is among the most popular choices. Alternative methods include pooling layers like GeM [45] or learned global aggregation, like the recent MixVPR [2]. In this paper, we propose optimal transport aggregation, setting a new state of the art in VPR.
@@
  • Secondly, we integrate the representational power of foundation models into VPR, using DINOv2 as the backbone for feature extraction. Unlike previous approaches that utilized DINOv2 in its pre-trained form, our method involves fine-tuning the model specifically for the task. This fine-tuning process converges extremely fast, in just four epochs, and allows DINOv2 to capture more relevant and distinctive features pertinent to place recognition tasks.
The fusion of these two novel components results in DINOv2 SALAD, which can be efficiently trained in less than one hour and sets unprecedented recall in VPR benchmarks, with 75.0% Recall@1 in MSLS Challenge and 76.0% in Nordland. All of this with a single-stage pipeline, without requiring expensive post-processing steps and with an inference speed of less than 3 ms per image.
@@ 3.1 Local Feature Extraction
DINOv2 adopts a ViT architecture that initially divides an input image $\mathbf{I}\in\mathbb{R}^{h\times w\times c}$ into $p\times p\times c$ patches, with $p=14$. These patches are sequentially projected with transformer blocks, resulting in the output tokens $\{\mathbf{t}_{1},\dots,\mathbf{t}_{n},\mathbf{t}_{n+1}\},\mathbf{t}_{i}\in\mathbb{R}^{d}$, where $n=hw/p^{2}$ is the number of input patches and there is an additional global token $\mathbf{t}_{n+1}$ that aggregates class information. Although the DINOv2’s authors reported that fine-tuning the model only brings dim improvements, we found that at least for VPR there are substantial gains in selectively unfreezing and training the last blocks of the encoder.
@@ 3.3 Aggregation
where $\textbf{t}_{n+1}$ is the global token from DINOv2. We then concatenate g with V flattened. Following NetVLAD, we do an L2 intra-normalization and an entire L2 normalization of this vector, which yields the final global descriptor.
@@ 4 Experiments
Desc. size Latency (ms) R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
NetVLAD [4] 32768 1.41 35.1 47.4 82.6 89.6 32.6 47.1 90.5 96.2 78.7 88.3
GeM [45] 1024 1.14 49.7 64.2 78.2 86.6 21.6 37.3 87.0 94.4 66.7 83.4
Conv-AP [1] 8192 1.22 54.2 66.6 83.1 90.3 42.7 58.9 92.9 97.7 79.2 88.6
CosPlace [5] 2048 2.59 67.2 78.0 87.4 93.0 44.2 59.7 92.1 97.5 80.1 89.6
MixVPR [2] 4096 1.37 64.0 75.9 88.0 92.7 58.4 74.6 94.6 98.3 85.2 92.1
EigenPlaces [6] 2048 2.65 67.4 77.1 89.3 93.7 54.4 68.8 94.1 98.0 69.9 82.9
DINOv2 SALAD 512 + 32 2.33 70.8 83.6 89.3 94.9 61.2 78.9 93.0 97.4 88.5 94.7
DINOv2 SALAD 2048 + 64 2.35 73.7 85.9 90.5 95.4 70.4 85.7 94.8 98.3 89.5 94.9
DINOv2 SALAD 8192 + 256 2.41 75.0 88.8 92.2 96.4 76.0 89.2 95.1 98.5 92.1 96.2
Table 1: Comparison against single-stage baselines. We compare DINOv2 SALAD against two popular baselines [4, 45] and the four baselines that show best results in recent literature [1, 5, 2, 6]. Our slim version already obtains state-of-the-art results in all metrics. Our full model outperforms all previous results by a significant margin. Note, in particular, the large improvement in the most challenging benchmarks, MSLS Challenge and NordLand. † We reproduced GeM results training during 80 epochs following MixVPR training pipeline.
Method Desc. size Memory (GB) Latency (ms) MSLS Challenge MSLS Val
@@
Table 2: Comparison against baselines with re-ranking. We compare our single-stage DINOv2 SALAD with methods that perform a re-ranking stage to improve performance. Without using re-ranking, our DINOv2 SALAD outperforms all other methods while being orders of magnitude faster and more memory-efficient. Latency metrics obtained from [65] using a RTX A5000. Latency for DINOv2 SALAD was computed using a RTX 3090. Memory footprint is calculated on the MSLS Val dataset, which includes around $18,000$ images.
Method Desc. size SF-XL Test v1 SF-XL Test v2
CosPlace [5] 2048 76.4 88.8
EigenPlaces [6] 2048 84.1 90.8
DINOv2 SALAD 8192 + 256 88.6 94.8
Table 3: Results on SF-XL. (R@1) Our DINOv2 SALAD achieves unprecedented results on SF-XL despite never seeing any single image of San Francisco during VPR finetuning.
To rigorously evaluate the effectiveness of our proposed contributions, we conducted exhaustive experiments following standard evaluation protocols.
@@ 4.1 Implementation Details
For the architecture, we opt for a pretrained DINOv2-B backbone, targeting a balance between computational efficiency and representational capacity. We only fine-tune the final 4 layers of the encoder, which significantly enhances the performance without markedly increasing training time. For the fully connected layers, the weights of the hidden layers $\mathbf{W}_{s_{1}}$, $\mathbf{W}_{f_{1}}$ and $\mathbf{W}_{g_{1}}$ have $512$ neurons and use ReLU for the activation function $\sigma$. To optimize feature handling, we employ a dimensionality reduction, compressing feature token dimensions from $d=768$ to $l=128$, and the global to 256. We use $m=64$ clusters, resulting in a global descriptor of size $128\times 64+256$. We also report results with smaller descriptors, with size $512+32$ ($m=15,\ l=32$), and $2048+64$ ($m=32,\ l=64$).
We train on GSV-Cities [1], a large dataset of urban locations collected from Google Street View. Given the impressive representation power of DINOv2, our pipeline achieves training convergence within just 4 epochs. Using a batch size of 60 places, each represented by 4 images, the training is completed in 30 minutes on a single NVIDIA RTX 3090. We use the multi-similarity loss [59] and AdamW [35] for the optimization, with an initial learning rate set to $6\mathrm{e}{-5}$. To ensure an effective learning rate, we linearly decay the initial rate at every iteration so at the end of the training is $20\%$ of the initial value. We use a dropout rate of $0.3$ on the score projection and dimensionality reduction neurons. As our model is agnostic to the image input size (as long as it can be divided in $14\times 14$ patches), we evaluate on images of size $322\times 322$ but train on $224\times 224$ to speedup training time.
To validate our experiments and select the hyperparameters, we monitored the recall in the Pittsburg30k-test [57]. We observed that, in the long run, most configurations perform similarly, but rapid convergence on a few epochs is more sensitive to the hyperparameters.
@@ 4.2 Results
We benchmarked our model against several single-stage baselines, namely NetVLAD [4] and GeM [45] as two representative tradicional baselines, and Conv-AP [1], CosPlace [5], MixVPR [2] and EigenPlaces [6] as the four most recent and best performing baselines in the literature. The evaluation spanned a diverse array of well-established datasets: MSLS Validation and Challenge [60], which are comprised of dashcam images; Pittsburgh250k-test [57], featuring urban scenarios; SPED [14], a collection from surveillance cameras; NordLand, notable for its seasonal variations from images captured from the front of a train traversing Norway; and SF-XL [5], a large urban dataset to evaluate VPR at scale. We use Recall@k (R@k) as the metric for all our experiments, as it is standard in related work. We use evaluation data and code from MixVPR [2], which considers retrieval as correct if an image at less than $25$ meters (or two frames for Nordland) from the query is among the top-k predicted candidates.
As shown in Table 1, our model outperforms all previous methods on all datasets and all metrics. Even the smaller $512+32$ version already surpasses previous models with bigger descriptors on most datasets. It is worth highlighting the metrics saturation observed in MSLS Val, Pitts250k-test and SPED, and on the other hand the challenging nature of MSLS Challenge and NordLand. The MSLS Challenge dataset, with its diversity, extensive size and closed labels, and NordLand, with its extreme sample similarity and seasonal shifts, emerge then as key benchmarks for assessing VPR performance. Although our DINOv2 SALAD shows a significant improvement on all benchmarks, it is precisely in MSLS Challenge and NordLand where we obtain the most substantial recall increases, with $+7.6\%,+11.7\%$ and $+17.6\%,+14.6\%$ for R@1, R@5 respectively over the second best. For SF-XL, as shown in Table 3, our method also achieves the best results to date. This is remarkable, considering that the previous state of the art was trained on this dataset, whereas our method never used any image of San Francisco when it was fine-tuned.
Method MSLS Challenge MSLS Val NordLand Pitts250k-test SPED
Desc. size R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
@@
In Table 2, we compare our DINOv2 SALAD method, which solely operates on a single retrieval stage, against the leading two-stage Visual Place Recognition (VPR) techniques. In this comparison, we include the best performing models in the literature, namely R2Former [65], TransVPR [58], and Patch-NetVLAD [23], which incorporate a re-ranking refinement. Note how our DINOv2 SALAD, despite being orders of magnitude faster and smaller in memory, significantly outperforms all these two-stage methods on all benchmarks. This finding not only highlights the efficiency of our model but also demonstrates the effectiveness of global retrieval using our novel SALAD aggregation. Additionally, considering our method’s reliance on local features, we believe that a re-ranking stage could also be applied, potentially increasing our recall metrics but at the price of a higher computational footprint.
@@ 4.3 Ablation Studies
Effect of DINOv2. We assess the impact of the DINOv2 backbone and our optimal transport aggregation SALAD separately. For this, we compare with the existing baselines of ResNet NetVLAD or AnyLoc, this last one applying a VLAD on top of a pretrained DINOv2 encoder. We integrate the DINOv2 backbone with various aggregation modules, obtaining a handful of performant techniques that improve their respective previous results. As shown in Table 4, all of these outperform the baselines, even though AnyLoc already uses DINOv2. This validates the DINOv2’s integration in end-to-end fine-tuning to refine its capabilities.
@@
Effect of SALAD. Our experiments in Table 4 show that aggregation also matters. Even the recent MixVPR aggregation coupled with DINOv2 does not match the performance of DINOv2 NetVLAD and DINOv2 SALAD. We believe that the DINOv2 backbone is especially suitable for local feature aggregation, as its features work remarkably well in dense visual perception tasks [41, 27, 62]. Although DINOv2 NetVLAD achieves comparable performance to SALAD, it employs a descriptor almost three times as big. Besides, the generalization performance of DINOv2 NetVLAD is limited, as observed in NordLand results. We attribute this to NetVLAD’s priors initialization with urban scenarios, which constrain the convergence of the system. In our experiments we also trained a slimmer DINOv2 NetVLAD version, whose features are dimensionally reduced as described in Section 3.3, targetting a final descriptor of roughly the same size as SALAD. In this fairer setup, DINOv2 SALAD clearly outperforms DINOv2 NetVLAD. We also evaluate SALAD on top of ResNet and ConvNext backbones, which improves over baseline ResNet NetVLAD but is significantly worse than using DINOv2. This indicates that SALAD is specially suited for high spatial resolution features, like the ones from DINOv2.
@@
Effect of hyperparameters. DINOv2 comes in different sizes that affect the number of parameters, inference speed, and representation capabilities. As shown in Table 5, more parameters do not always result in better performance. Excessively big models might be harder to train or prone to overfitting the training set. From these results, we chose the DINOv2-B backbone, which exhibits a great balance between performance and size and speed. Regarding descriptor size, we observed (Table 1) that changing $m$ and $l$ allows to get slimmer versions with competitive performance. For the number of blocks to train, as shown in Table 6, fine-tuning two or four block report the best results without significant computation overhead.
Model Dim. size # Params. Latency (ms) MSLS Val R@1
S 384 21M 1.30 90.5
B 768 86M 2.41 92.2
L 1024 300M 7.82 92.6
G 1536 1100M 24.93 91.7
Table 5: DINOv2 configurations and performances.
@@
Figure 3: Heatmap of local features importance. Left images show the original pictures, their right counterparts represent the weights not assigned to the ‘dustbin’. Note how the network learns to discard uninformative regions like skies, roads or dynamic objects, and instead focus on distinctive patterns in buildings and vegetation. We attribute its focus on distant buildings to their invariance to viewpoint change.
@@
Effect of SALAD components. In Table 7, we show how different components of our SALAD pipeline affect the final performance. Both the global token, which appends global information not captured in local features, and the dustbin, which helps in distilling the aggregated features, contribute to the performance of SALAD. We also trained a model using a dual-softmax [46] to solve the optimal transport assignment, following LoFTR and Gluestick [52, 42]. Although dual-softmax achieves only slightly worse performance, the Sinkhorn Algorithm is theoretically sound and provides a better acronym to our method.
@@
Table 6: Fine-tuning different number of DINOv2 blocks.
Method MSLS Val
R@1 R@5 R@10
DINOv2 SALAD w/o dustbin 91.4 95.8 96.2
DINOv2 SALAD w/o global token 91.8 96.0 96.2
DINOv2 SALAD (Dual Softmax) 91.9 95.7 96.5
DINOv2 SALAD 92.2 96.4 97.0
Table 7: Ablation study of the SALAD components.